OpenStudy (anonymous):

how long will it take to double an investment at 6.25% interest if compounded monthly

OpenStudy (anonymous):

6.25 percent interest compounded monthly means that you every month you multiply your investment by 1 + 0.0625/12 = 193/192. Now since you're doubling, you have 2 = (193/192)^t where t is the number of months. Obviously you'll want to round up if you want to actually double since you don't get the interest until the end of the month.